Determine the fourth roots of 16 in the form x+iy.where x and y are not trigonometric functions

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)
Answer by stanbon(75887)

Determine the fourth roots of 16 in the form x+iy.where x and y are not trigonometric functions
-------------
16 = 16(cos(0+2npi) + isin(0+2npi)
-----
16^(1/4) = 2(cos](0+2npi)/4]+ isin[(0+2npi)/4]
----
Let n = 0 to get 16^(1/4) = 2(cos(0) + isin(0) = 2 + 0i
Let n = 1 to get 2[cos(pi/2)+isin(pi/2)] = 2(0+i) = 0+2i
Let n = 2 to get 2(cos(pi) + isin(pi)] = 2(-1+0i) = -2+0i
Let n = 3 to get 2[cos(3pi/2) + isin(3pi/2)] = 2(0-i) = 0-2i
======================
Cheers,
Stan H.
----------------------